OpenStudy (anonymous):

help please! A. Line m contains the points, A(–2, 6) and B(4, 8), while line n contains the C(8, 12) and D(x, 24). Given m and n are perpendicular lines, solve for the value of x. Given m and n are parallel lines, solve for the value of x. In your final answer, include all formulas and calculations necessary to solve for x.

OpenStudy (swamijatin2):

You know two point, the intercepts. you can use those to find the slope pretty easy. Then put it in the right form.

OpenStudy (swamijatin2):

Which should tell you the slope right off.

OpenStudy (swamijatin2):

and for y=mx+b form you need the y intercept for b... well... you have it. So you have b.

OpenStudy (swamijatin2):

Well, what is the y intercept? Remember, in y=mx+b the b is the y intercept and the m is the slope. If you have those two things you can then put them in the formula and move it around to get standard form. solve y=−x+4 for standard form. Remember that earlier you agreed on ax+by=c as the standard form. So you need the x and y on the left with only a number on the right. =) So one last step and you are done. m+n−i−42=0 To move the 42, I would need to add to both sides: m+n−i−42+42=0+42 m+n−i=42 To move the i, becvause it is -i, I need to add it: m+n−i+i=42+i m+n=42+i To move the n, because it is +, I must subtract it. m+n−n=42+i−n m=42+i−n So you must do the opposite of what sign it is. You have a −x so you must add it. y=(−1)x+4 means the -1 is multiplying x. So it becomes: y=−1x+4 y=−x+4

OpenStudy (swamijatin2):

Well, I hope you can see the process. It was really just the same as if you were given 2 points at random. But with that graph, the two points were easy ones to work with.

OpenStudy (anonymous):

that doesnt help much, i have to solve for X in both cases.

OpenStudy (anonymous):

@swamijatin2